[figure: 9.7 cm]

HYPERFINE FIELDS AT THE ^{119}Sn NUCLEAR SITES EMBEDDED IN Cr LAYERS OF Fe/Cr MULTILAYERS

K. Mibu, M. Almokhtar M., S. Tanaka, and T. Shinjo

Institute for Chemical Research, Kyoto University, Uji, Kyoto-fu 611-0011, Japan

In order to obtain information about the local magnetic properties of Cr layers in Fe/Cr multilayer systems, the magnetic hyperfine fields induced at the ^{119}Sn nuclear sites of epitaxial Fe/Cr/Sn/Cr multilayers with monatomic Sn layers were studied through^{119}Sn conversion electron Mössbauer spectroscopy. The results were compared with those for Cr/Sn multilayers [1]. The hyperfine fields at the Sn nuclear sites of Fe/Cr/Sn/Cr multilayers at room temperature decrease from 9 to 2 T as the Cr layer thickness decreases from 40 to 5 Å, whereas those for Cr/Sn multilayers increase a little from 11 to 12 T in the same thickness region. The difference in magnetic hyperfine fields in both systems is thought to be due to an influence of Fe on the magnetic properties of Cr [2]. Possible magnetic structures of the Cr layers are inferred from the size and direction of the magnetic hyperfine field transferred at the Sn nuclear sites.

[1] K. Mibu et al., J. Phys. Soc. Jpn. 67 (1998) 2633.

[2] K. Mibu et al., J. Magn. Magn. Mater. (1999), to be published.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : oral

Corresponding author : K. Mibu

Address for correspondence : Institute for Chemical Research, Kyoto University Uji, Kyoto-fu 611-0011, Japan

Email address : mibu@scl.kyoto-u.ac.jp